"Two sides of a square were increased by 6 cm. If the other two opposite sides were decreased by 6 cm, by how much does the area change?"
Answer by mananth(16946)

let side be x
Area = x^2
New Length = (x+6)
New width = (x-6)
Area = (x+6)(x-6)
= x^2-36
x^2-x^2+36 is the difference
the difference is 36 cm^2
m.ananth@hotmail.ca